OpenStudy (anonymous):

find y'' of xy + tan^-1 (xy) = 4 a. -(2y/x^2) b. -(2x/y^2) c. 2y/x^2 d. 2x/y^2

OpenStudy (anonymous):

\[y'=-\frac{y}{x}\] is a start

OpenStudy (anonymous):

then

OpenStudy (anonymous):

i get letter A when i do that but i dont trust always my answer

OpenStudy (ash2326):

We have \[xy+\tan ^{-1} (xy)=4\] Let's differentiate this we get we know \[\frac{d}{dx} \tan^{-1} x=\frac{1}{1+x^2}\] \[y+ x\frac{dy}{dx}+ \frac{1}{1+x^2y^2} (y+x\frac{dy}{dx})=0\] Let's take the common term out \[(y+x\frac{dy}{dx})(1+\frac{1}{1+x^2y^2})=0\] we get \[(y+x\frac{dy}{dx})(1+x^2y^2+1)=0\] or \[(y+x\frac{dy}{dx})(2+x^2y^2)=0\] 2+x^y^ can't be zero, so \[y+x\frac{dy}{dx}=0\] or \[ \frac{dy}{dx}=-\frac{y}{x}\] we have \[y+x\frac{dy}{dx}=0\] Let's differentiate this \[\frac{dy}{dx}+ \frac{dy}{dx}+x\frac{d^2y}{dx^2}=0\] now we get \[2\frac{dy}{dx}+x\frac{d^2y}{dx^2}=0\] or \[\frac{d^2y}{dx^2}=-2\frac{dy}{xdx}\] we have \( \frac{dy}{dx}=-\frac{y}{x}\) so \[\frac{d^2y}{dx^2}=-2\frac{-y}{x\times x}\] we get finally \[\frac{d^2y}{dx^2}=\frac{2y}{x^2}\]

OpenStudy (anonymous):

omg thats a big help

OpenStudy (ash2326):

:D :D

OpenStudy (anonymous):

can i learn this thing in 1 day or not?

OpenStudy (ash2326):

You want to learn differentiation?

OpenStudy (anonymous):

yes i do but how many days do i need to learn this thing? For example the topic right now on school is about derivative of trigonometric function.

OpenStudy (ash2326):

Just go along with the class. Learn the topic the same day when it's taught in School. It'll be a slow process but you'll learn it beautifully:)

OpenStudy (anonymous):

maybe i do that after class i will try my best.